tne Hoss. A ssume ha dliag omls cans or pre ssine Sofce Determine Capproneler tha farce cach nembar of agenals can suppart eithar a tensile. ek com
media%2Fa86%2Fa866ca4c-3efd-4a1e-9abc-54
3 K 20 ーナー-22H
Answer #1

Let Va and Vd be vertical reactions at Support A and D.

so, Va + Vd = 4+8+8+4 = 24

Now moment about A = 0

so, Vd*60 - 4*60 - 8*40 - 8*20 - 3*20 = 0

i.e, Vd = 780/60 = 13 k (upwards)

so, Va = 24 - 13 = 11 k (upwards)

As Hinge support is only at A, so it will resist horizontal support.

so, Ha = 3 k (towards left)

Now to calculate force in each element we will take each joint one by one.

ah IK Fhb FA K 4S CTeasin) 3 -91 8.7 11-3 Va-

Member Force (K) Nature (Compression/Tension)
AH 11 C
AB 3 T
AG 0
HG 8.7 C
BG 8 C
BH 9.9 T
GC 0
BF 11.31 T
GF 8.7 C
BC 8 C
FC 16.7 C
FD 12.3 T
CE 0
FE 8.7 C
CD 8 C
DE 4 C